OpenStudy (anonymous):

Math Analysis Ch. 10.4 Solve the equation for 0

sam (.sam.):

Use trig identity

OpenStudy (anonymous):

can you list the steps, i really don't get it :\

OpenStudy (anonymous):

would i - the 3sin x over to the other side?

sam (.sam.):

cos(2x) = cos^2(x) - sin^2(x) = 2 cos^2(x) - 1 = 1 - 2 sin^2(x)

sam (.sam.):

3sin x = 1+ cos 2x 3sin x = 1+ ( 1 - 2 sin^2(x)) 2 sin^2(x)+3sin x-2=0 (sinx+2)(2sinx-1)=0 sinx=-2 ,sinx=1/2

OpenStudy (anonymous):

okay, ty :) How about number 2?

sam (.sam.):

Change the secx to 1/(cosx) then change the cos2x to 2 cos^2(x) - 1

OpenStudy (anonymous):

then i'd minus 2 cos^2(x) -1 to the other side?

sam (.sam.):

make it like a quadratic and solve by factoring

OpenStudy (anonymous):

so i would make 1 have a denominator of cosx so i can be able to add them together?

OpenStudy (anonymous):

like this...? 0=1/cosx - 2cos^2x + 1 0=1/cosx - 2cos^2x +cos/cosx 0=1+cosx/cosx - 2cos^2x

sam (.sam.):

cos2x=secx 2cos^2x-1=1/cosx 2cos^3x-cosx=1 cosx-1=0 cosx=1

OpenStudy (anonymous):

how did you get 2cos^3x?

OpenStudy (anonymous):

wait nvm

OpenStudy (anonymous):

so how did you get.. cosx-1=0 cosx=1

sam (.sam.):

factor the cubic equation

OpenStudy (anonymous):

so when i get .. 2cos^3-cosx=1 i would minus 1 to the other side again?

sam (.sam.):

yes

OpenStudy (anonymous):

why..

sam (.sam.):

2cos^3-cosx=1 2cos^3-cosx-1=0 factor

OpenStudy (anonymous):

ok so i got.. cosx(2cos^2x-1)=1

OpenStudy (anonymous):

cosx=1 2cos^2x-1=1 +1+1 2cos^2x/2=2/2 square root of(cos^2x)=1 cosx=1..?

OpenStudy (anonymous):

so both answers would be cosx=1 o: RIGHT? :p

sam (.sam.):

dont do 2cos^2x-1=1

sam (.sam.):

use cosx=1 only

OpenStudy (anonymous):

why not, that's what i got when i factored...

sam (.sam.):

you cant factor when there's 1 at the right side, the only thing to factor is right side is zero 2cos^3-cosx-1=0 so you'll get cosx-1=0 cosx=1

OpenStudy (anonymous):

yah, and so i factored 2cos^3x-cosx-1=0 and i got.. cosx(2cos^2x-1)=1

OpenStudy (anonymous):

i mean 2cos^3x-cosx=1

OpenStudy (anonymous):

like you said an hour ago.

sam (.sam.):

2cos^(3)x-cosx-1=0 Factoring you get (cosx-1)(2cos^(2)x+2cosx+1)=0 (cosx-1)=0, (2cos^(2)x+2cosx+1)=0 From cosx-1=0 cosx=1 \[x=\cos^{-1}(1)\] x=0 ---------------------------------------------- For the quadratic, use quadratic formula 2cos^(2)x+2cosx+1=0 cosx=(-b+-sqrt(b^(2)-4ac))/(2a) cosx=(-2+-sqrt((2)^(2)-4(2)(1)))/(2(2)) cosx=(-2+-2i)/2(2) cosx=(-2+-2i)/4 cosx=(-1+i)/2 cosx=(-1-i)/2 cosx=(-1+i)/(2),(-1-i)/2 ------------------------------------ cosx=(-1+i)/(2)_cosx=(-1-i)/(2) cosx=(-1+i)/2 cosx=(-1-i)/2 You don't get any real solution for the quadratic, the only real solution is x=0,180,360

OpenStudy (anonymous):

why would x=0,180,and360?

OpenStudy (anonymous):

cosx=1 is only on 0 lolz.

sam (.sam.):

0 and 360

sam (.sam.):